OpenStudy (anonymous):

Write the equation for the line tangent to the circle x2 + y2 = 29 at the point (2, 5).

OpenStudy (anonymous):

Im so confused can anyone explain this to me? PLEASE

OpenStudy (anonymous):

eq. of tangent is xx1+yy1=a^2 2x+5y=29

OpenStudy (anonymous):

do you know this formula ? or you have to use dy/dx at (2,5) this is slope of tangent at (2,5) then eq. of tangent is y-y1=(dy/dx)(x-x1)

OpenStudy (anonymous):

\[2x+2y \frac{ dy }{ dx }=0\]

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=-\frac{ 2x }{ 2y }=-\frac{ x }{ y }\] at (2,5)\[\frac{ dy }{dx }=-\frac{ 2 }{ 5 }\] slope of tangent=-2/5 \[eq.~ of~ tangent~ is~ y-y1=m(x-x1)\] \[y-5=-\frac{ 2 }{ 5 }\left( x-2 \right)\] solve it. which will be same as listed above.